suppose that f is a linear function with slop 4 and that f(3) = 5. find the equasion for f

Answer by nerdybill(7384)

From:suppose that f is a linear function with slope 4 and that f(3) = 5. find the equation for f
.
When you see:
f(3) = 5
this basically gives you a "point" on the line at (3,5)
.
That, along with the slope (4) can be used in the "point-slope" form:
y – y1 = m(x – x1)
.
Plugging in our give values:
y – y1 = m(x – x1)
y – 5 = 4(x – 3)
y – 5 = 4x – 12
y = 4x -7 (This is the equation for f)